# Small Language Models: Survey, Measurements, and Insights

### v1

@@ 1 Introduction
The evolution of language models is diverging. On the one hand, in pursuit of artificial general intelligence following the scaling law, larger and larger language models (LLM) have been born in datacenters that host hundreds of thousands of GPUs [34, 71]. The ultimate goal of this path is to demonstrate that machines can solve the most challenging language tasks and greatly advance human civilization (e.g., science). On the other hand, small language models (SLM) are proposed for resource-efficient deployment, especially on devices such as desktops, smartphones, and even wearables. A vision for SLMs is to democratize machine intelligence, making it accessible and affordable to anyone, anywhere, at any time – just like the human brain that everyone possesses.
@@
Both LLM and SLM are important in reshaping our daily lives, yet the latter receives significantly less attention in academia. There has been very limited literature that understands SLM’s capability [37, 59, 79] or their runtime cost on devices [41, 36, 69], often with limited scale or depth. In the real world, however, SLMs have already been integrated into commercial off-the-shelf (COTS) devices on a massive scale [74]. For instance, the latest Google/Samsung smartphones have built-in LLM services (Gemini Nano), allowing third-party mobile apps to leverage LLM capabilities through prompts and LoRA modules [1]. The most recent iOS system on iPhones and iPads also includes an on-device foundation model, deeply integrated with the operating system for better performance and privacy [2].
@@
This work presents the first comprehensive survey of SLMs, thoroughly discussing their capabilities, runtime costs, and innovations in recent years. The scope of this survey is limited to those language models with 100M–5B parameters in decoder-only transformer architecture***The definition of “small” is subjective and relative. It could also drift over time, considering that device memory is increasing over time and can host larger “small language models” in the future. We set 5B as the upper limit for the size of SLMs, since as of Seq. 2024, 7B LLMs are still mostly deployed in the cloud, at least in Sep. 2024., which cover the device range from low-end IoT/wearable gadgets like smartwatches to high-end mobile devices such as smartphones and tablets. In total, we collected 57 popular SLMs and fully benchmarked their capabilities (commonsense reasoning, math, in-context learning, long-context retrieval, etc.) and on-device runtime costs (prefill and decode speed, memory footprint, energy consumption, etc.).
@@
With such in-depth investigation, we obtain valuable insights that help answer critical questions such as: “What is the evolving path of SLMs?” “What datasets or training strategies are more likely to produce a highly capable SLM?” “How different SLM architecture (e.g., depth, width, atten type) and the deployment environments (quantization algorithms, hardware type, etc) impact runtime performance?” We expect the work to reveal an all-sided landscape of SLM and benefit the research community, including those working on the algorithm, model, system, and hardware levels.
Overall, we make the following contributions in this work.
@@ 2.3 Training Datasets
We also studied the open-sourced pre-training datasets quality based on the performance of SLMs trained on them. We classified the SLMs in the past three years into groups of less than 0.5B, 1B, 2B, and 3B according to the number of parameters, and sorted them based on the average accuracy (the average of the accuracy of the two types of tasks, Commonsense reasoning/understanding and Problem solving, as will be shown in later section) of the four groups of models to explore how to select datasets to improve the average accuracy. The results are shown in Table 2. We notice that two recently released datasets, DCLM and FineWeb-Edu, show superior performance compared to others. One common feature of the two datasets is the adoption of model-based data filtering. In addition, coding data is often included in the SLM pretraining datasets such as StarCoder, even though coding ability is not a focus of SLMs deployed on devices. This is likely attributed to the common belief that coding data can help improve the model reasoning ability [78].
@@
The number of parameters in SLM models and the amount of data used for training (the number of tokens) are closely related, with the Chinchilla law [32] suggesting that the optimal ratio between the number of model parameters and training tokens should be around 20 (e.g., a 1B model with 20B tokens). We have statistically analyzed the number of training tokens used by SLMs under 4B parameters from 2022 to 2024, as shown in Figure 4(a). Generally, the larger the model, the greater the number of tokens used for training, and more recent models tend to have more training tokens. A notable observation is that SLMs are trained on much large number of tokens (typically over 1.5T) than what is suggested by the Chinchilla law, regardless of their parameter sizes. In some cases, smaller SLMs are trained on even more data than the larger SLMs (e.g., Qwen2-0.5B with 12T tokens compared to Qwen2-1.5B with 7T tokens). This indicates that these SLMs are significantly ”over-trained”. The rationale behind this approach is to deploy powerful SLMs on resource-constrained devices by using more training-time FLOPs.
@@
Insights: We make two key observations in SLM training datasets. Data quality is crucial to SLM capability, which receives increasing attentions in recent SLM research. The importance of data quality to the final SLM capability typically outweighs the data quantity and model architecture configurations. A notable trend of dataset research is using model-based filtering, which result in two state-of-the-art open-sourced pre-training datasets: FineWeb-Edu (1.3T/5.4T) [55] and DCLM-baseline (4T) [39]. SLMs trained on these two datasets have achieved competitive performance to those trained on closed datasets, which have significantly advanced the SLM research reproducibility. Recent SLMs are trained over large amount of tokens (typically $>$1.5T), disregarding their parameter sizes. In some cases, smaller SLMs are trained over even more data (e.g., Qwen2-0.5B at 12T tokens but Qwen2-1.5B at 7T tokens). It also means those SLMs are significantly “over-trained”, as compared to the Chinchilla law [32] that estimates the parameter-token ratio to be around only 20 (e.g., 1B model with 20B tokens). The incentive of such “over-training” action is to deploy powerful SLMs on resource-constrained devices through investing more training-time compute resources.
@@ 3.2 Overall Capabilities
(a) Commensense reasoning/understanding.
(b) Problem solving.
(c) Math.
@@
Insights: We draw four key insights from the development of SLMs: From 2022 to 2024, SLMs exhibited significant performance improvements across various language tasks, outpacing the improvements of the LLaMA-7B series (1/2/3/3.1 versions). This paints a promising picture for SLMs’ potential to solve a range of downstream tasks on devices. The Phi family consistently achieves state-of-the-art performance across most tasks. In particular, Phi-3-mini achieves the highest accuracy as of September 2024, rivaling LLaMA 3.1 8B. While much of its superior performance may be due to careful data engineering by the Microsoft team, part of it may also be attributed to instructive tuning and potential overfitting to specific datasets [77]. Although larger parameter counts generally lead to better performance, exceptions such as Qwen 2 1.5B demonstrate that smaller models can still excel in specific tasks. SLMs trained on open-source datasets are closing the gap with their closed-source counterparts in commonsense tasks. However, the gap remains significant in tasks requiring complex reasoning or logic. This underscores the need for improved datasets focused on mathematical reasoning to address this disparity.
@@ 4 SLM Runtime Cost
We evaluated 20 models on two types of edge devices: the Jetson Orin Module, commonly used in edge AI devices such as drones and small robots, and a smartphone, which people rely on in their daily lives. The detailed specifications are shown in Table 3. All experiments on the Jetson used its GPU, while those on the smartphone were performed using its CPU. To eliminate the impact of inference engine implementations, we carry out all experiments using llama.cpp, a widely recognized open-source inference engine.
Device Name Hardware Type Specifications
Jetson Orin NX 16GB GPU 1024-core NVIDIA Ampere architecture GPU with 32 tensor cores, 16G DRAM
MEIZU 18Pro CPU Snapdragon 888, 8G RAM
Table 3: Experiment devices.
We primarily recorded metrics as model parameter, latency during the prefill and decode phases, and runtime memory usage. Due to variations in how each model officially reports its parameter counts, we relied on the parameter values obtained from llama.cpp. Inference is divided into two phases: prefilling and decoding. During the prefill phase, the input prompt is processed to generate a KV Cache, where multiple tokens in the prompt can be computed in parallel. For the prefill phase, we focused on first token latency, which represents the time it takes to process all tokens in the prompt. The decode phase, also known as the autoregressive phase, generates one token at a time, incorporating it into the KV Cache. Simultaneously, this token is used in predicting the next token. For the decode phase, we measured latency per token. We set a standard prompt length of 50 and a token generation length of 50 unless specified otherwise. Tests are conducted at 10-second intervals to mitigate potential thermal throttling issues. To measure larger models, we applied 4-bit quantization to all models before conducting experiments in all sections except §$\S$ 4.2.1. Therefore, the latency and memory usage reported are models after 4-bit quantization.
@@ 4.2.1 Impact of Quantization
Insights: We have two key insights about the impact of quantization methods for latency with different prompt length and output token length. The benefits of quantization during the decode stage are greater than those in the prefill stage. On mobile devices, quantization mainly reduces memory access overhead. Since the decode stage is more bandwidth-bound, it gains more from quantization compared to the compute-bound prefill stage. More regular quantization precision leads to better performance. Although 3-bit quantization offers a higher model compression rate, 4-bit quantization performs better in both the prefill and decode stages. The inferior performance of 3-bit quantization is due to its irregular bit-width, which lacks hardware optimization support and incurs additional overhead from data alignment and padding. As a result, despite its lower compression rate, 4-bit quantization is more efficient. Similarly, irregular 5-bit and 6-bit quantization result in inference latency that is comparable to, or even higher than 8-bit quantization, despite offering higher compression rates.
@@ 4.2.2 Impact of Hardware
Figure 9: Latency under different hardware.
@@
Insights: We have two key insights about the impact of hardware for latency with different prompt length and output token length. GPU shows an even greater advantage over the CPU during the prefill phase. The prefill phase involves parallel processing of tokens within the prompt, whereas the decode phase generates each token sequentially. Therefore, the prefill phase has a higher degree of parallelism, making it more suitable for GPUs, which have more parallel computing units. The Jetson demonstrates better performance stability compared to the smartphone. Due to its relatively simple hardware structure, which facilitates better heat dissipation, the Jetson maintains more stable latency during lengthy inference tasks.

### v2

@@ 1 Introduction
The evolution of language models is diverging. On one hand, in the pursuit of artificial general intelligence (AGI) and following the scaling law, increasingly large language models (LLM) have been born in datacenters that host hundreds of thousands of GPUs [40, 94]. The aim of this path is to demonstrate that machines can solve the most challenging language tasks, with the ultimate goal of advancing human civilization by pushing the boundaries of science and technology. On the other hand, there is a growing focus on small language models (SLM), designed for resource-efficient deployment on personal devices such as desktops, smartphones, and even wearables. The vision behind SLMs is to democratize access to machine intelligence, making it both accessible and affordable to people everywhere. This approach seeks to make intelligence ubiquitous and practical, available to anyone, anywhere, at any time – much like the human brain, which everyone possesses.
@@
Both LLM and SLM are important in reshaping our daily lives, yet the latter receives significantly less attention in academia. There has been very limited literature exploring SLM capabilities [44, 74, 108] or their runtime cost on devices [50, 43, 92], often with limited scale or depth. In the real world, however, SLMs have already been integrated into commercial off-the-shelf (COTS) devices on a massive scale [99, 25]. For instance, the latest Google/Samsung smartphones have built-in LLM services (Gemini Nano), allowing third-party mobile apps to leverage LLM capabilities through prompts and LoRA modules [1]. The most recent iOS system on iPhones and iPads also includes an on-device foundation model, deeply integrated with the operating system for better performance and privacy [2]. Beyond resource-constrained scenarios, SLMs also excel superior performance in certain domain-specific tasks [46].
@@
This work presents the first comprehensive survey of SLMs, thoroughly discussing their capabilities, runtime costs, and innovations in recent years. The scope of this survey is limited to those language models with 100M–5B parameters in decoder-only transformer architecture***The definition of “small” could drift over time, considering that device memory is increasing over time and can host larger “small language models” in the future. We set 5B as the upper limit for the size of SLMs, since as of Seq. 2024, 7B LLMs are still mostly deployed in the cloud., which covers the range of devices from low-end IoT/wearable gadgets like smartwatches to high-end mobile devices such as smartphones and tablets. In total, we collected 57 popular SLMs and fully benchmarked their capabilities (commonsense reasoning, math, in-context learning, long-context retrieval, etc.) and on-device runtime costs (prefill and decode speed, memory footprint, energy consumption, etc.).
@@
With the insights from comprehensive benchmarking and thorough investigation, We try to answer the following questions concerning SLMs: “What is the evolving path of SLMs?” “What datasets or training strategies are more likely to produce a highly capable SLM?” “How different SLM architecture (e.g., depth, width, atten type) and the deployment environments (quantization algorithms, hardware type, etc) impact runtime performance?” We expect the work to reveal an all-sided landscape of SLM and benefit the research community, including those working on the algorithm, model, system, and hardware levels.
In summary, we make the following contributions in this work.
@@ 1.7B MHA 24 2048 32
Princeton [Uncaptioned image] Sheared-LLaMA [72] 1.3B 2023.11 MHA 24 2048 16 SiLU 32k 4096
2.7B 2023.11 MHA 32 2560 20 SiLU 32k 4096
Xiaohongshu [Uncaptioned image] MiniMA [88] 3B 2023.11 UKN 24 3072 24 SiLU 49k 4096
MiniMA2 [89] 1B 2024.07 18 2304 18 SiLU 49k 4096
Nvidia [Uncaptioned image] Minitron [66] 4B 2024.07 GQA 32 3072 24 ReLU2 256k 4096
M.A.P. [Uncaptioned image] CT-LLM [55] 2B 2024.04 MHA 32 2048 16 SiLU 125k 4096
AMD [Uncaptioned image] AMD-Llama [9] 135M 2024.08 MHA 12 768 12 SiLU 32k 2048
@@ 2.3 Training Datasets
Comparing the quality of pre-training datasets. We also studied the open-sourced pre-training datasets quality based on the performance of SLMs trained on them. We classified the SLMs in the past three years into groups of less than 0.5B, 1B, 2B, and 3B according to the number of parameters, and sorted them based on the average accuracy (the average of the accuracy of the two types of tasks, Commonsense reasoning/understanding and Problem solving, as will be shown in later section) of the four groups of models to explore how to select datasets to improve the average accuracy. The results are shown in Table 2. We notice that two recently released datasets, DCLM and FineWeb-Edu, show superior performance compared to others. One common feature of the two datasets is the adoption of model-based data filtering. In addition, coding data is often included in the SLM pretraining datasets such as StarCoder, even though coding ability is not a focus of SLMs deployed on devices. This is likely attributed to the common belief that coding data can help improve the model reasoning ability [106].
@@
The number of parameters in SLM models and the amount of data used for training (the number of tokens) are closely related, with the Chinchilla law [37] suggesting that the optimal ratio between the number of model parameters and training tokens should be around 20 (e.g., a 1B model with 20B tokens). We have statistically analyzed the number of training tokens used by SLMs under 4B parameters from 2022 to 2024, as shown in Figure 4(a). Generally, the larger the model, the greater the number of tokens used for training, and more recent models tend to have more training tokens. A notable observation is that SLMs are trained on much large number of tokens (typically over 1.5T) than what is suggested by the Chinchilla law, regardless of their parameter sizes. In some cases, smaller SLMs are trained on even more data than the larger SLMs (e.g., Qwen2-0.5B with 12T tokens compared to Qwen2-1.5B with 7T tokens). This indicates that these SLMs are significantly ”over-trained”. The rationale behind this approach is to deploy powerful SLMs on resource-constrained devices by using more training-time FLOPs. Though, SLMs are known to have performance saturation issue with over-training [31].
@@
Insights: We make two key observations in SLM training datasets. Data quality is crucial to SLM capability, which receives increasing attentions in recent SLM research. The importance of data quality to the final SLM capability typically outweighs the data quantity and model architecture configurations. A notable trend of dataset research is using model-based filtering, which result in two state-of-the-art open-sourced pre-training datasets: FineWeb-Edu (1.3T/5.4T) [69] and DCLM-baseline (4T) [47]. SLMs trained on these two datasets have achieved competitive performance to those trained on closed datasets, which have significantly advanced the SLM research reproducibility.
Recent SLMs are trained over large amount of tokens (typically $>$1.5T), disregarding their parameter sizes. In some cases, smaller SLMs are trained over even more data (e.g., Qwen2-0.5B at 12T tokens but Qwen2-1.5B at 7T tokens). It also means those SLMs are significantly “over-trained”, as compared to the Chinchilla law [37] that estimates the parameter-token ratio to be around only 20 (e.g., 1B model with 20B tokens). The incentive of such “over-training” action is to deploy powerful SLMs on resource-constrained devices through investing more training-time compute resources.
@@ 3.2 Overall Capabilities
Insights: We draw four key insights from the development of SLMs: From 2022 to 2024, SLMs exhibited significant performance improvements across various language tasks, outpacing the improvements of the LLaMA-7B series (1/2/3/3.1 versions). This paints a promising picture for SLMs’ potential to solve a range of downstream tasks on devices.
The Phi family consistently achieves state-of-the-art performance across most tasks. In particular, Phi-3-mini achieves the highest accuracy as of September 2024, rivaling LLaMA 3.1 8B. While much of its superior performance may be due to careful data engineering by the Microsoft team, part of it may also be attributed to instructive tuning and potential overfitting to specific datasets [103]. Although larger parameter counts generally lead to better performance, exceptions such as Qwen 2 1.5B demonstrate that smaller models can still excel in specific tasks. SLMs trained on open-source datasets are closing the gap with their closed-source counterparts in commonsense tasks. However, the gap remains significant in tasks requiring complex reasoning or logic. This underscores the need for improved datasets focused on mathematical reasoning to address this disparity.
@@ 3.3 In-context Learning Capabilities
3.4 Long Context Capabilities
(a) Llama3.2-1B
(b) Llama3.2-3B
(c) Qwen1.5-0.5B
(d) Qwen1.5-1.8B
(e) Qwen2-0.5B
(f) Qwen2-1.5B
(g) Qwen2.5-0.5B
(h) Qwen2.5-3B
Figure 7: Needle In A Haystack
We used Needle-In-A-Haystack provided by OpenCompass to explore long-context capabilities of SLMs. The tasks included Single-Needle Retrieval, Multi-Needle Retrieval, and Multi-Needle Reasoning. The scores in Figure 7 are the average of these three tasks. Different models showed large variations in performance. Small models, such as Qwen1.5-0.5B and Qwen2-0.5B, performed less effectively. Qwen1.5-0.5B achieved an average accuracy of 22.13%. Qwen2-0.5B performed slightly better, reaching 43.84%. Qwen1.5-0.5B handled shorter contexts (9k-17k) relatively well. However, its accuracy dropped sharply with longer contexts. This was especially true for middle inserted text (Depth Percent from 20% to 70%). Larger models performed much better. Llama 3.2-3B had an average accuracy of 57.81%. It worked well with shorter contexts but struggled with deeper insertions when contexts exceeded 25k tokens. Qwen2.5-3B achieved an average accuracy of 91.71%. It maintained nearly perfect accuracy across all context lengths and insertion positions. This result highlights its strong ability to handle long contexts and complex scenarios.
Insights: We draw two key insights from the long context capacity of SLMs: Larger parameters are crucial for long-context capabilities. Small models, such as Qwen1.5-0.5B and Qwen2-0.5B, perform adequately on short-context tasks but experience a significant drop in recognition accuracy as the context length increases. In contrast, larger models, such as Qwen2.5-3B, excel with outstanding performance, maintaining near-perfect accuracy across all context lengths and insertion positions. ”Lost in the Middle” also occurs in small models. Compared to deep or front insertions, the accuracy of middle-position text (Depth Percent 20%-70%) is significantly lower.
@@ 4 SLM Runtime Cost
We evaluated 20 models on two types of edge devices: the Jetson Orin Module, commonly used in edge AI devices such as drones and small robots, and smartphones, which people rely on in their daily lives. The detailed specifications are shown in Table 3. All experiments on the Jetson used its GPU, while those on the smartphones were performed using its CPU. To eliminate the impact of inference engine implementations, we carry out all experiments using llama.cpp, a widely recognized open-source inference engine.
Device Name Specifications Release Time
Jetson Orin NX 16GB 1024-core NVIDIA Ampere architecture GPU with 32 tensor cores, 16G DRAM Feb. 2023
Pixel 7Pro GoogleTensor G2,12G RAM Oct. 2022
Xiaomi 12S Snapdragon 8Gen1+ ,12G RAM Jul. 2022
MEIZU 18Pro Snapdragon 888,8G RAM Mar. 2021
Table 3: Testing devices.
We primarily recorded metrics as model parameter, latency during the prefill and decode phases, and runtime memory usage. Due to variations in how each model officially reports its parameter counts, we relied on the parameter values obtained from llama.cpp. Inference is divided into two phases: prefilling and decoding. During the prefill phase, the input prompt is processed to generate a KV Cache, where multiple tokens in the prompt can be computed in parallel. For the prefill phase, we focused on first token latency, which represents the time it takes to process all tokens in the prompt. The decode phase, also known as the autoregressive phase, generates one token at a time, incorporating it into the KV Cache. Simultaneously, this token is used in predicting the next token. For the decode phase, we measured latency per token. We set a standard prompt length of 50 and a token generation length of 50 unless specified otherwise. Tests are conducted at 10-second intervals to mitigate potential thermal throttling issues. To measure larger models, we applied 4-bit quantization to all models before conducting experiments in all sections except §$\S$ 4.2.1. Therefore, the latency and memory usage reported are models after 4-bit quantization. Further optimization in recent literature is thoroughly discussed in §$\S$ 4.4.
@@ 4.2.1 Impact of Quantization
Insights: We have two key insights about the impact of quantization methods for latency with different prompt length and output token length. The benefits of quantization during the decode stage are greater than those in the prefill stage. On mobile devices, quantization mainly reduces memory access overhead. Since the decode stage is more bandwidth-bound, it gains more from quantization compared to the compute-bound prefill stage. The benefits of quantization during the prefill stage decrease with prompt length increasing. Quantization compresses weights and kv cache. All tokens share the weight file so the benefit for each token decreases when there are more tokens in prompt. However, decode stage has more memory operation in kv cache so it still is benefited with generation tokens increasing. More regular quantization precision leads to better performance. Although 3-bit quantization offers a higher model compression rate, 4-bit quantization performs better in both the prefill and decode stages. The inferior performance of 3-bit quantization is due to its irregular bit-width, which lacks hardware optimization support and incurs additional overhead from data alignment and padding. As a result, despite its lower compression rate, 4-bit quantization is more efficient. Similarly, irregular 5-bit and 6-bit quantization result in inference latency that is comparable to, or even higher than 8-bit quantization, despite offering higher compression rates.
@@ 4.2.2 Impact of Hardware
Figure 10: Latency on GPU and CPU.
(a) Prefill
(b) Decode
Figure 11: Latency on different smartphones.
@@
We also tested the prefill and decode times of Qwen1.5-1.8B on three smartphones. To minimize the impact of power consumption, a 60-second interval was set between each test. As shown in the figures, the latency for prefill and decode on the three smartphones increases linearly with the number of tokens. The Xiaomi 12S performed the best with the lowest latency, highlighting the efficiency of the Snapdragon 8Gen1+ chip. The Pixel 7 Pro followed, delivering competitive performance, while the MEIZU 18 Pro had the highest latency due to its older Snapdragon 888 chip and lower memory configuration.
Insights: We have two key insights about the impact of hardware for latency with different prompt length and output token length. GPU shows an even greater advantage over the CPU during the prefill phase. The prefill phase involves parallel processing of tokens within the prompt, whereas the decode phase generates each token sequentially. Therefore, the prefill phase has a higher degree of parallelism, making it more suitable for GPUs, which have more parallel computing units. The Jetson demonstrates better performance stability compared to the smartphone. Due to its relatively simple hardware structure, which facilitates better heat dissipation, the Jetson maintains more stable latency during lengthy inference tasks. The development of System on a Chip (SoC) generations effectively improves inference efficiency.
@@ 4.3.2 Memory Breakdown
4.4 Optimizations for On-device Deployment
Since SLMs are deployed on resource-constrained devices, numerous approaches have been developed to optimize their performance in terms of latency, memory, and other overheads. These methods can be categorized into two types: Online and Offline. Online optimization focuses on enhancing SLM performance during runtime, with key techniques including hardware-aware optimizations, model collaboration and quantization. In contrast, Offline optimization targets SLM improvements prior to deployment, including optimizing datasets, model architecture and knowledge distillation.
4.4.1 Online Optimization
Hardware-aware optimization requires fully leveraging the heterogeneous computing capabilities of edge devices. This involves optimizing task scheduling across CPUs, GPUs, and NPUs, as well as effectively utilizing multiple storage hierarchies, such as cache, DRAM, and flash. Earlier frameworks were primarily designed for CPU-GPU orchestration. EdgeNN [102] optimizes inference on IoT devices with a CPU-GPU architecture by improving the use of unified memory and enhancing task scheduling between the CPU and GPU, bringing an average of 3.97×, 3.12×, and 8.80× speedups to inference on the CPU of the integrated device, inference on a mobile phone CPU, and inference on an edge CPU device, respectively. Transformer-Lite [48] is a framework for efficiently deploying large language models on mobile phone GPUs by optimizing memory usage and computational efficiency. It utilizes techniques like model pruning, quantization, and layer fusion to reduce memory footprint and computational load, achieving up to 5.7x faster inference compared to CPU-based methods while maintaining model accuracy. PowerInfer [78] introduces an efficient framework for serving large language models (LLMs) on consumer-grade GPUs. It optimizes inference by partitioning the model into cold-activated neurons and hot-activated neurons, and assigns the heavy computation tasks to the GPU while offloading lighter tasks to the CPU. This results in up to 2.4x faster performance, reduced memory usage, and lower power consumption, making LLM inference feasible on consumer-grade hardware.
Recent systems like PowerInfer-2 [95] and mllm-NPU [91] are specifically designed for off-the-shelf modern smartphones, where the AI computing power resides in the NPU. Consequently, both frameworks focus on optimizing the CPU-NPU architecture. PowerInfer-2, as an extension of the original PowerInfer, adopts a neuron cluster approach, where matrix operations are decomposed into clusters of neurons. Based on the activation sparsity, it schedules the computation on either the NPU or CPU. This efficient allocation of resources results in up to a 27.8x speed improvement compared to existing state-of-the-art frameworks. Notably, PowerInfer-2 can serve a 47B parameter LLM at an impressive rate of 11.68 tokens per second, enabling fast and efficient LLM inference on smartphones without significant accuracy degradation. mllm-NPU optimizes at three levels—prompt, tensor, and block and achieves up to 22.4x faster prefill speed, 30.7x energy savings, and up to 32.8x speedup in end-to-end real-world applications, with the ability to process over 1,000 tokens per second for Qwen1.5-1.8B. RIPPLE [85] proposes the concept of Neuron Co-Activation, where neurons frequently activated together are linked to facilitate continuous read access and optimize data transfer efficiency. It achieves up to 5.93× improvements in I/O latency compared to the state-of-the-art.
For the memory optimization, LLMS [97] has proposed an effective solution. To reduce the context switching overhead when running LLMs under tight device memory constraints, LLMS employs fine-grained memory management by decoupling the memory management of the application and LLM. It utilizes globally optimized key-value (KV) cache compression and swapping techniques to efficiently manage memory resources. LLMS achieves up to 100x faster performance in context-switching compared to existing solutions on various edge devices.
The deployment of mixture-of-experts (MoE) models on edge devices also faces challenges in terms of huge memory usage. EdgeMoE [96] is the first on-device inference engine for MoE-based large language models, optimizing memory and computation by storing non-expert weights in device memory and expert weights in external storage, fetched only when activated. It further improves efficiency through expert-wise bitwidth adaptation and advanced expert management techniques. It reduces memory footprint by 1.05x to 1.18x compared to traditional models that store the entire model in memory. Additionally, EdgeMoE demonstrates 1.5x to 2.5x faster inference on Jetson TX2 (GPU) and Raspberry Pi 4B (CPU), with some tasks showing up to 10x speedup due to optimized memory and computation management. MoE Cache-Conditional [75] also chooses to keep non-expert weights resident in DRAM, while expert weights are dynamically scheduled from flash storage back to DRAM. It present on-device results demonstrating 2× speedups on mobile devices.
Speculative sampling is commonly used on the cloud to reduce the computational overhead of LLMs. This idea of model collaboration can similarly be applied effectively on SLMs. LLMCad [90] is an efficient on-device inference engine for generative NLP tasks on mobile devices. By combining a smaller memory-resident LLM for token generation and a larger LLM for token validation, LLMCad achieves up to 9.3x faster performance compared to existing engines on IoT devices and smartphones, respectively, without comprising accuracy.
Besides validating the outputs of smaller models with large models, models of different scales can also be applied to different tasks. ELMS [98] proposes a one-time neuron reordering technique for efficient sub-model generation and a dual-head compact language model for elastic mobile LLM services. ELMS improves accuracy by up to 16.83% on end-to-end traces and speeds up inference time by up to 5× with less than 1% switching overhead, while maintaining comparable memory usage and utilizing fewer than 100 GPU hours for model preparation.
4.4.2 Offline Optimization
Datasets play a crucial role in the accuracy of SLMs. Recent studies have focused on enhancing model performance by optimizing data utilization or generating specific types of data for training. TinyStories [26] is a synthetic dataset designed for training small language models. Trained on those child-friendly vocabulary, models as small as 28M parameters can produce coherent stories comparable to GPT-2-XL (1.5B parameters). Models trained on TinyStories require less than one GPU-day for training while exhibiting emergent reasoning and language capabilities. AS-ES learning [87] is a novel data-efficient training paradigm for improving the reasoning performance of small models by segmenting chain-of-thought data into Abstractive Segments and Extractive Segments. This method enables iterative generation without requiring additional data or modifications to the model architecture. On math word problems, AS-ES learning improves accuracy by up to 15.28%, while reducing BLEU scores, suggesting a trade-off that favors logical reasoning over memorization. For PET scan summarization tasks, it achieves up to 30.6% higher ROUGE-L scores, highlighting its effectiveness across diverse reasoning-intensive tasks. Self-AMPLIFY [14] improves small language models’ reasoning ability by generating rationales using post hoc explanation methods, eliminating the need for auxiliary models or human annotations
Model architecture optimization involves enhancing the model’s efficiency and performance by adjusting or redesigning its architecture. OnceNAS [107] is a neural architecture search method designed for on-device inference on edge devices with constrained resources. By leveraging a continuous latent space for architecture representation and integrating parameter count, latency, and accuracy as optimization objectives, OnceNAS achieves a 10.49× size reduction and a 5.45× speedup compared to baseline methods. Evaluations across various edge devices, including Raspberry Pi and FPGA, demonstrate superior efficiency and generalizability, making OnceNAS highly suitable for edge intelligence applications like autonomous driving and smart healthcare. Weight-Inherited Distillation (WID) [86] proposes a task-agnostic BERT compression technique that eliminates alignment losses by directly inheriting weights through structural re-parameterization with row and column compactors. WID achieves up to 49.2% parameter reduction while retaining 98.9% of BERT-base performance on GLUE benchmarks and significantly outperforms baselines like TinyBERT and MiniLM in both accuracy and efficiency, with training time reduced by over 50%.
Knowledge Distillation transfers knowledge from larger models to smaller ones through knowledge distillation, thereby improving the performance of smaller models, particularly in reasoning tasks. DISTILLM [41] proposes an efficient knowledge distillation framework, leveraging a novel skew Kullback-Leibler divergence loss and an adaptive off-policy approach to address training-inference mismatches and enhance computational efficiency. Compared to existing KD methods, DISTILLM achieves up to 4.3× faster training speeds while maintaining or improving performance on OpenLLaMA2-3B.  [54] explores transferring reasoning capabilities from large language models to smaller ones via chain-of-thought knowledge distillation. By fine-tuning smaller models like T5 on CoT outputs generated by LLMs such as PaLM 540B and GPT-3 175B, the authors achieve substantial improvements in reasoning tasks. For instance, T5 XXL’s accuracy on the GSM8K dataset increased from 8.11% to 21.99% when fine-tuned on CoT data generated by PaLM 540B, highlighting a 170% improvement. The approach also demonstrates robustness across arithmetic, commonsense, and symbolic reasoning tasks.
Quantization is another effective method for model compression and reducing operational overhead, but there are few methods specifically designed for edge devices. Activation-aware Weight Quantization (AWQ) [51] enables low-bit weight-only quantization for LLMs on edge devices. AWQ protects 1% of the most critical weights using per-channel scaling informed by activation patterns, achieving a 3-4× speedup over FP16 implementations and reducing memory usage by 4×. AWQ is paired with TinyChat, a framework that further optimizes on-device inference, achieving up to 3.3× acceleration on GPUs.  [84] compares FP8 and INT8 formats for efficient on-device deep learning inference. It demonstrates that while FP8 may offer slightly improved accuracy for certain outlier-dominated tasks, INT8 achieves superior overall efficiency, reducing hardware costs by up to 50% and energy usage by 53%. For post-training quantization and quantization-aware training, INT8 consistently outperforms FP8 in networks with well-behaved distributions, proving more robust and computationally efficient across a range of tasks.